For a continuous random variable that has a normal distribution with mean of 24 and a standard deviation of 6, find the area under the normal curve from 𝑥=18 and 𝑥=21.


1
Expert's answer
2022-03-27T16:51:53-0400

We have a normal distribution, "\mu=24, \sigma=6."

Let's convert it to the standard normal distribution, "z=\cfrac{x-\mu}{\sigma};"

"z_1=\cfrac{18-24}{6}=-1, \\z_2=\cfrac{21-24}{6}=-0.5,"


"P(18<X<21)=P(-1<Z<-0.5)=\\=P(Z<-0.5)-P(Z<-1)="

"=0.30854-0.15866=0.14988" ​(from z-table).